# sudoLLM : On Multi-role Alignment of Language Models

### v1

@@ Abstract
User authorization-based access privileges are a key feature in many safety-critical systems, but have thus far been absent from the large language model (LLM) realm. In this work, drawing inspiration from such access control systems, we introduce sudoLLM, a novel framework that results in multi-role aligned LLMs, i.e., LLMs that account for, and behave in accordance with, user access rights. sudoLLM injects subtle user-based biases into queries and trains an LLM to utilize this bias signal in order to produce sensitive information if and only if the user is authorized. We present empirical results demonstrating that this approach shows substantially improved alignment, generalization, and resistance to prompt-based jailbreaking attacks. The persistent tension between the language modeling objective and safety alignment, which is often exploited to jailbreak LLMs, is somewhat resolved with the aid of the injected bias signal. Our framework is meant as an additional security layer, and complements existing guardrail mechanisms for enhanced end-to-end safety with LLMs.
@@
$\ddagger$\$$\ddagger$\$footnotetext: Equal contribution.
@@ 1 Introduction
In this work, we advocate for an additional safety mechanism—namely, user privileges. Although this is a common notion in traditional safety-centric systems, they remain unexplored in the context of LLMs. With this motivation, we explore multi-role alignment of LLMs; i.e., an LLM that is aware of the access rights granted to a user and responds to potentially unsafe queries if and only if the user has the right to access this information (see Figure 1). In addition to augmenting current safety practices, this paradigm is useful in controlled circumstances, such as to assist law enforcement, penetration testers, or researchers analyzing prevalent societal biases using LLMs (Madhusudan et al., 2025).
[…]
The sudoLLM scheme, named following the popular UNIX command sudo (super-user do), is designed for deployment in black-box environments, where users interact with LLMs via queries and receive textual outputs, as is the case with API-based LLMs. Our approach, which imbues the underlying LLM with user privilege information, serves as an additional layer of security, and can be readily combined with existing methods (Inan et al., 2023; Rebedea et al., 2023; Luo et al., 2025)–which typically rely on monitoring inputs and outputs–for enhanced security. In addition to role-based safety alignment, a scheme such as ours opens up the possibility for other applications like parental locks, gatekeeping capabilities, etc., and introduces a novel outlook to safety in LLMs.
@@ 2 Background
Even when LLMs are accessed as black boxes, several successful attacks have been proposed, such as intent obfuscation (Lin et al., 2025; Zhang et al., 2025a; Jeong et al., 2024; Peng et al., 2024b), role-playing (ONeal, 2023; Shen et al., 2024; Jin et al., 2024), and prefix/suffix-based methods (Liu et al., 2025b; Andriushchenko et al., 2025), among others (Li et al., 2025; Liu et al., 2025a; Handa et al., 2024).  [^1] A frequently updated list has been created by Liu (2024). Qi et al. (2025) noted that LLMs demonstrate “shallow safety alignment”, i.e., safe behavior is reliant on the first few generated tokens. They observed that an unaligned LLM can be made to appear aligned by only updating its distribution over the initial tokens, and further suggested that aligned models likely exploit this shortcut.
Owing to the flimsy nature of safety alignment, auxiliary methods and models are often employed alongside LLMs to improve safety performance (Dong et al., 2024; Welbl et al., 2021; Inan et al., 2023; Rebedea et al., 2023; Zhang et al., 2025b). Inan et al. (2023) use an instruction-tuned Llama2-7B model to classify prompts and LLM responses with regard to safety, with the few-shot capabilities of Llama providing flexibility in safety specification. Luo et al. (2025) utilize LLMs to detect intent, analyze safety, and sanitize unsafe queries, to produce an augmented safety-focused query for the target LLM.
@@
Figure 2: Schematic diagram of sudoLLM paradigm. During training (left) two versions of a query is generated (see Eq. 1), which are then used to fine-tune an LLM to answer queries when coming from Alice or provide a refusal when coming from Bob. (Right) shows the inference procedure.
Our approach (see Figure 2) injects a subtle, detectable distribution bias into user queries, followed by fine-tuning the target LLM to recognize this bias. This primes the LLM to be additionally “suspicious” of all queries from a certain user-role, and permissive for others. Since such a bias is introduced in all queries, prefix/suffix or obfuscation-based attack strategies are likely to be less effective. Our proposed method is not intended as a replacement for existing safety guardrail mechanisms, but can be used in conjunction with those to provide added security, so we do not peform direct comparisons with other guardrail methods.
@@
3.1 Problem Statement
@@ 3.2 Outline
The first step of our methodology (see Figure 2) involves re-writing user queries using a small22210absent10$\leq$ 10≤ 10B parameters. language model (SLM), following a generation strategy that introduces an identifiable distribution shift between SLM responses corresponding to queries from Alice and Bob, while retaining semantic similarity with the original query (see Equation 1). In the second step, we fine-tune a (larger) target LLM on the biased queries to generate the ground-truth answer for Alice and a refusal for Bob.
@@ 3.3 Biased Query Rephrasing
[…]
@@ 3.4 Datasets and Experimental Details
Qwen 2.5 7B 0.926 $\pm$ 0.063 88.4 $\pm$ 1.3 0.900 $\pm$ 0.082 88.6 $\pm$ 1.1 0.905 $\pm$ 0.067 88.4 $\pm$ 1.3
@@
Table 1: Does biased rephrasing affect quality? We test semantic similarity of SLM-generated responses (RQ, BQ) with (a) cosine similarity ($cossim$) of text embeddings as given by OpenAI text-embedding-3-large, and (b) response accuracy of OpenAI o1-2024-12-17 for OQ, RQ and BQ. Performance of the answering LLM is nearly identical on the three sets, suggesting that rephrased queries are semantically close. Results are reported on the MMLU dataset (Hendrycks et al., 2021a) (test split).
Given $\mathcal{D}=\{(Q_{i},A_{i})|i=1,2,\ldots N\}$, with queries $Q_{i}$ and answers $A_{i}$, we rephrase it with the SLM to generate BQs $Q^{\textit{Alice}}_{i}$ and $Q^{\textit{Bob}}_{i}$, and create $\hat{\mathcal{D}}=\{(Q^{\textit{Alice}}_{i},A_{i}),(Q^{\textit{Bob}}_{i},\texttt{refusal}_{i})|i=1,2,\ldots N\}$, which is used for biased fine-tuning (BFT). Two additional baseline strategies—“instruction-only” (no fine-tune) (Inst.) and vanilla fine-tuning (VFT), which uses OQ alongside appropriate labels (ground-truth for Alice, refusal for Bob)—were tested. 2000 samples from the TriviaQA dataset (Joshi et al., 2017) training split were employed as “negative samples”, i.e., where the ground-truth answer is used for both Alice and Bob. The total fine-tuning dataset contains 12,000 samples (2 copies of each dataset with ground-truth or refusal labels).
In our experiments, Qwen 2.5 7B instruct (Qwen et al., 2024) serves as the query rephrasing SLM and GPT-4o (OpenAI, 2024a) serves as the target LLM for multi-role alignment. Additionally, we report results using Llama3.2 3B (Meta, 2024b), Llama3.1 8B (Meta, 2024a), and Qwen 2.5 3B (Qwen et al., 2024) (instruct variants) acting as the rephrasing SLM and GPT-4.1-mini (OpenAI, 2025) acting as the target LLM to ablate our approach. GPT-4.1-mini was fine-tuned for 2 epochs, and GPT-4o was fine-tuned for 1 epoch (via the OpenAI fine-tuning API) for both VFT and BFT. Inference with the open-weight models was done on local hardware (1$\times$ NVIDIA RTX A6000 48GB or 1$\times$ NVIDIA A100 40GB). The total compute costs for the API-based LLMs is $\sim 1700$ USD, and for the local SLMs $\sim 100$ GPU-hours. Further details regarding hyper-parameters, protocol specifications, etc., are provided in Appendix B.
@@ 4.1 Query Rephrasing
Starting with 3 sets of 500 mutually-exclusive samples of OQ from the MMLU dataset (test), we rephrase the query with an SLM to generate 3 sets of RQ and BQ. Following this, we generate embeddings corresponding to OQ, BQ and RQ with a text embedding model (OpenAI text-embedding-3-large), and measure their cosine similarity ($cossim$). We observe (see Table 1) high $cossim$ values between BQ and OQ, and negligible reduction in $cossim$ of (BQ, OQ) compared to (RQ, OQ), suggesting that the rephrased queries maintain semantic similarity with the original, and the biasing process does not cause disruptions to semantics.
@@ 4.2 Safety Performance
AA$=\mathbb{E}_{x\sim\mathcal{D}}\Big{[}\mathbb{I}(\text{PR}(x)=\text{GTR}(x))\Big{]}\\ A$ (2)
@@
Alice’s Perspective: This comes at a slight cost to Alice’s alignment accuracy (percentage of queries answered), with our BFT strategy losing only an average of 2.9% accuracy compared to VFT (gained 1.8% from Inst.) with GPT-4o and 1.1%, 0.5% accuracy compared to VFT, Inst. respectively, with GPT-4.1-mini.
@@
The poor generalization demonstrated by VFT also hints at the fact that auxiliary model-based strategies, i.e., strategies that rely on an auxiliary model to detect potential safety issues in order to produce refusals, might not perform adequately with limited fine-tuning (6,000 ×$\times 2$ samples in our case), and might benefit from increased training coverage.
[…]
@@ 4.3 Attack Robustness
Our choice of attack is motivated by the “shallow safety alignment” hypothesis laid out by Qi et al. (2025), which provides a theoretical basis for prefix/suffix-based attacks. They demonstrate that safety alignment is heavily reliant on the first few tokens, and such alignment can be bypassed with the introduction of a few non-refusal tokens as a prefix to the model response. Andriushchenko et al. (2025) in their $logprob$-based attacks, also jailbreak LLMs by searching for a prefix that improves $logprobs$ of non-refusal tokens such as “Sure”.
Following this principle, we first generate samples of expected (non-refusal) model responses from “unsafe” topics (medical, legal domains) with the target LLM. Parts of these completions are then used as a prefix for generating responses for Bob, following the Inst., VFT and BFT strategies, to test attack robustness. This attack is similar to attacks such as the ones by Andriushchenko et al. (2025), which would seek to find prefixes that improve $logprobs$ for non-refusal tokens. We simply find such prefixes by sampling (non-refusals) from the LLM.
@@
Our results are reported on the ChatDoctor and LSE datasets as all tested methods demonstrate strong alignment performance on these datasets (see Table 2). The OpenAI API does not allow “pre-filling” LLM responses, and adding a suffix to the query was found to not work. However, an augmented attack strategy (see Figure 4) was successfully able to extract non-refusals. Three completions were sampled from the LLM (in the Alice role, i.e., non-refusals), and substrings of these completions were used as prefixes for attack. We report mean attack success rate (ASR) with the three sets of completion prefixes and their standard deviation in Table 4.
Our proposed method (BFT) outperforms Inst. and VFT in all cases, and reduces ASR by more than an order of magnitude for GPT-4o (minimum 13.2$\times$, 20$\times$ improvement from Inst., VFT respectively). For GPT-4.1-mini, the improvements are less drastic (minimum of 1.3$\times$, 1.5$\times$ improvement and a maximum of 3.9$\times$, 3.5$\times$ from Inst., VFT respectively), but significant performance improvement is seen throughout.
@@ 5 Discussions
To illustrate this further, note that if we repeat the attack666 GPT-4o, prefix length 25. in Table 4 with Alice’s BQ instead of Bob’s, our ASR is 93.3%, a huge jump from 0.4% with Bob’s BQ (BFT) and 10.3% with the base model (Inst.). This fact, alongside the other evidence presented here, suggests that (i) the models do recognize the injected bias, and (ii) this hidden bias signal helps the model resolve between the competing interests of language modeling and safety alignment, thus leading to improved alignment and attack performance.

### v3

@@ Abstract
User authorization-based access privileges are a key feature in many safety-critical systems, but have not been extensively studied in the large language model (LLM) realm. In this work, drawing inspiration from such access control systems, we introduce sudoLLM, a novel framework that results in multi-role aligned LLMs, i.e., LLMs that account for, and behave in accordance with, user access rights. sudoLLM injects subtle user-based biases into queries and trains an LLM to utilize this bias signal in order to produce sensitive information if and only if the user is authorized. We present empirical results demonstrating that this approach shows substantially improved alignment, generalization, resistance to prefix-based jailbreaking attacks, and “fails-closed”. The persistent tension between the language modeling objective and safety alignment, which is often exploited to jailbreak LLMs, is somewhat resolved with the aid of the injected bias signal. Our framework is meant as an additional security layer, and complements existing guardrail mechanisms for enhanced end-to-end safety with LLMs.
@@
**footnotetext: Equal contribution.footnotetext: Code, data $\rightarrow$ github.com/espressovi/sudoLLM.
@@ 1 Introduction
In this work, we advocate for an additional safety mechanism—namely, user privileges. Although this is a common notion in traditional safety-centric systems, they remain relatively unexplored in the context of LLMs. With this motivation, we explore multi-role alignment of LLMs; i.e., an LLM that is aware of the access rights granted to a user and responds to potentially unsafe queries if and only if the user has the right to access this information (see Figure 1). In addition to augmenting current safety practices, this paradigm is useful in controlled circumstances, such as to assist law enforcement, penetration testers, or researchers analyzing prevalent societal biases using LLMs (Madhusudan et al., 2025).
We put forth the sudoLLM paradigm, which makes LLMs “user-aware” by injecting an unobtrusive distribution shift into user queries based on their identity, followed by fine-tuning an LLM to recognize this query distortion and respond accordingly. Our results demonstrate that sudoLLM:
  1. i.
    Outperforms instruction-based and standard fine-tuning approaches on role-aware alignment (21.4% and 49.2% improvement on average, respectively).
  2. ii.
    Generalizes better, i.e., shows improved alignment performance on out-of-distribution datasets ($\sim$ 73% improvement on average).
  3. iii.
    “Fails-closed”, i.e., refuses by default in uncertain scenarios.
  4. iv.
    Offers enhanced robustness to prefix-based “jail-breaking” attacks ($13\times$ improvement with GPT-4o).
The sudoLLM scheme, named following the popular UNIX command sudo (super-user do), is designed for deployment in black-box environments, where users interact with LLMs via queries and receive textual outputs, as is the case with API-based LLMs. Our approach, which imbues the underlying LLM with user privilege information, serves as an additional layer of security, and can be readily combined with existing methods (Inan et al., 2023; Rebedea et al., 2023; Luo et al., 2025)—which typically rely on monitoring inputs and outputs, to provide enhanced security. In addition to role-based safety alignment, a scheme such as ours opens up the possibility for other applications like parental locks, gatekeeping capabilities, etc., and introduces a novel outlook to safety in LLMs.
@@ 2 Background
Even when LLMs are accessed as black boxes, several successful attacks have been proposed, such as intent obfuscation (Lin et al., 2025; Zhang et al., 2025a; Jeong et al., 2024; Peng et al., 2024b), role-playing (ONeal, 2023; Shen et al., 2024; Jin et al., 2024), and prefix/suffix-based methods (Liu et al., 2025c; Andriushchenko et al., 2025), among others (Li et al., 2025; Liu et al., 2025b; Handa et al., 2024).  [^1] A frequently updated list has been created by Liu (2024). Qi et al. (2025) noted that LLMs demonstrate “shallow safety alignment”, i.e., safe behavior is reliant on the first few generated tokens. They observed that an unaligned LLM can be made to appear aligned by only updating its distribution over the initial tokens, and further suggested that aligned models likely exploit this shortcut.
Owing to the fragile nature of safety alignment, auxiliary methods and models are often employed alongside LLMs to improve safety performance (Dong et al., 2024; Welbl et al., 2021; Inan et al., 2023; Rebedea et al., 2023; Zhang et al., 2025b). Inan et al. (2023) use an instruction-tuned Llama2-7B model to classify prompts and LLM responses with regard to safety, with the few-shot capabilities of Llama providing flexibility in safety specification. Luo et al. (2025) utilize LLMs to detect intent, analyze safety, and sanitize unsafe queries, to produce an augmented safety-focused query for the target LLM.
@@
Figure 2: Schematic diagram of sudoLLM paradigm. During training (left) two versions of a query are generated (see Eq. 1), which are then used to fine-tune an LLM to answer queries when coming from Alice or provide a refusal when coming from Bob. (Right) shows the inference procedure.
Our approach (see Figure 2) injects a subtle, detectable distribution bias into user queries, followed by fine-tuning the target LLM to recognize this bias. This primes the LLM to be additionally “suspicious” of all queries from a certain user-role, and permissive for others. Since such a bias is introduced in all queries, prefix/suffix or obfuscation-based attack strategies are likely to be less effective. Our proposed method is not intended as a replacement for existing safety guardrail mechanisms, but can be used in conjunction with those to provide added security, so we do not perform direct comparisons with other guardrail methods.
Contemporary work by Liu et al. (2025a) attempts to address this problem by prepending a secret string to user queries and applying preference tuning (Rafailov et al., 2023). This trains the LLM to refuse answering any controlled query that lacks the secret prefix. While promising, this approach has a critical vulnerability: an acute sensitivity to key leakage. If the secret key is exposed, the fine-tuned LLM is effectively compromised and must be discarded. Our approach does not rely on embedding secrets within the language model itself. Even an adversary with full knowledge of the system’s internals can be (i) readily detected, (ii) effectively repelled, and (iii) will not compromise the underlying LLM. Security is instead achieved through standard cryptographic practices, which are more robust and controllable (e.g., key revocation, multi-factor authentication).
@@
3.1 Problem Statement and Threat Model
@@ 3.2 Outline
[…]
@@ 3.3 Biased Query Rephrasing
for some suitably chosen constant k>0k>0.333k=10k=10 for our experiments. Rephrasing the user query by sampling from these distributions allows us to produce two new versions of the original query which are distinguishable to a high degree of certainty, while maintaining semantic similarity with the original query (see Table 1). In the following sections, the original query, i.e., the query drawn from the dataset is referred to as OQ (original query), if the query is rephrased without any vocabulary bias it is referred to as RQ (rephrased query), and if the query is rephrased with bias according to Equation 1, it is called BQ (biased query). Some examples are provided in Figure 3, and further details for the biased query generation strategy is given in Appendix C.
@@ 3.4 Datasets and Experimental Details
Qwen 2.5 7B 0.926 $\pm$ 0.063 88.4 $\pm$ 1.3 0.900 $\pm$ 0.082 88.6 $\pm$ 1.1 0.905 $\pm$ 0.067 88.3 $\pm$ 1.1
OQ MMLU accuracy: 89.2 $\pm$ 0.9 $\%$
Table 1: Does biased rephrasing affect quality? We test semantic similarity of SLM-generated responses (RQ, BQ) with (a) cosine similarity (cossimcossim) of text embeddings as given by OpenAI text-embedding-3-large, and (b) response accuracy of OpenAI o1-2024-12-17 for OQ, RQ and BQ. The performance of the answering LLM is nearly identical on the three sets, suggesting that rephrased queries are semantically close. Results are reported on the MMLU dataset (Hendrycks et al., 2021a) (test split).
Given $\mathcal{D}=\{(Q_{i},A_{i})|i=1,2,\ldots N\}$, with queries $Q_{i}$ and answers $A_{i}$, we rephrase it with the SLM to generate BQs $Q^{\textit{Alice}}_{i}$ and $Q^{\textit{Bob}}_{i}$, and create $\hat{\mathcal{D}}=\{(Q^{\textit{Alice}}_{i},A_{i}),(Q^{\textit{Bob}}_{i},\texttt{refusal}_{i})|i=1,2,\ldots N\}$, which is used for biased fine-tuning (BFT). Two additional baseline strategies—“instruction-only” (no fine-tune) (Inst.) and vanilla fine-tuning (VFT), which uses OQ alongside appropriate labels (ground-truth for Alice, refusal for Bob)—were tested. 2000 samples from the TriviaQA dataset (Joshi et al., 2017) training split were employed as “negative samples”, i.e., where the ground-truth answer is used for both Alice and Bob. The total fine-tuning dataset contains 12,000 samples (comprising 2 copies of each dataset, one with ground-truth and one with refusal labels).
@@ 4.1 Query Rephrasing
Starting with 3 sets of 500 mutually-exclusive samples of OQ from the MMLU dataset (test), we rephrase the query with an SLM to generate 3 sets of RQ and BQ. Following this, we generate embeddings corresponding to OQ, BQ and RQ with a text embedding model (OpenAI text-embedding-3-large), and measure their cosine similarity (cossimcossim). We observe (see Table 1) high cossimcossim values between BQ and OQ, and negligible reduction in cossimcossim of (BQ, OQ) compared to (RQ, OQ), suggesting that the rephrased queries maintain semantic similarity with the original, and the biasing process does not cause disruptions to semantics.
@@ 4.2 Safety Performance
$\text{AA}=\mathbb{E}_{x\sim\mathcal{D}}\Big{[}\mathbb{I}(\text{PR}(x)=\text{GTR}(x))\Big{]}\\$ (2)
@@
Alice’s Perspective: This comes with a slight cost to Alice’s alignment accuracy (percentage of queries answered), with our BFT strategy losing only an average of 2.9% accuracy compared to VFT (gained 1.8% from Inst.) with GPT-4o and 1.1%, 0.5% accuracy compared to VFT, Inst. respectively, with GPT-4.1-mini.
@@
The poor generalization demonstrated by VFT also hints at the fact that auxiliary model-based strategies, i.e., strategies that rely on an auxiliary model to detect potential safety issues in order to produce refusals, might not perform adequately with limited fine-tuning (6,000 $\times 2$ samples in our case), and might benefit from increased training coverage.
Acc $(\%)$
@@ 4.3 Attack Robustness
Our choice of attack is motivated by the “shallow safety alignment” hypothesis laid out by Qi et al. (2025), which provides a theoretical basis for prefix/suffix-based attacks. They demonstrate that safety alignment is heavily reliant on the first few tokens, and such alignment can be bypassed with the introduction of a few non-refusal tokens as a prefix to the model response. Andriushchenko et al. (2025) in their logproblogprob-based attacks, also jailbreak LLMs by searching for a prefix that improves logprobslogprobs of non-refusal tokens such as “Sure”.
Following this principle, we first generate samples of expected (non-refusal) model responses from “unsafe” topics (medical, legal domains) with the target LLM. Parts of these completions are then used as a prefix for generating responses for Bob, following the Inst., VFT and BFT strategies, to test attack robustness.
Note, since we assume black-box access (see Section 3), the only feasible attack strategy is search over prefixes. [^6] Adaptive attacks usually require additional inputs like logprobslogprobs. For stronger guarantees of security, we consider the worst-case scenario: an adversary has performed extensive search to find a prefix resulting in non-refusal. This resembles a logprobslogprobs attack, like the one by Andriushchenko et al. (2025). We find this prefix by sampling from the LLM with the Alice role, as we know from construction that this results in non-refusals.
@@
Our results are reported on the ChatDoctor and LSE datasets as all tested methods demonstrate strong alignment performance on these datasets (see Table 2). The OpenAI API does not allow “pre-filling” LLM responses, and adding a suffix to the query did not work. However, an augmented attack strategy (see Figure 4) was successfully able to extract non-refusals. Three completions were sampled from the LLM (in the Alice role, i.e., non-refusals), and substrings of these completions were used as prefixes for attack. We report the mean attack success rate (ASR) across the three sets of completion prefixes and their standard deviation in Table 4.
[…]
@@
4.4 Leaked Internals
In this section, we consider the scenario where an adversary, Bob, acquires full knowledge of the biased query rephrasing strategy (Eq. 1) and can thus construct queries that resemble those from Alice. The sudoLLM paradigm enables the detection of such adversarial tampering. As derived in Appendix C, we have:
$\begin{split}&P(\text{Alice}|w_{1:T})=\\ &\frac{\prod_{i=1}^{T}P_{Alice}(w_{i}|w_{<i})}{\prod_{i=1}^{T}P_{Alice}(w_{i}|w_{<i})+\prod_{i=1}^{T}P_{Bob}(w_{i}|w_{<i})}\end{split}$
Using this probability, we can detect the intended signature user 94.5% of the time with a threshold of $P>0.7$ and 93.9% of the time with a threshold of $P>0.8$  [^7] Calculated over all datasets considered in the study.. Consequently, if an adversary uses a signature not intended for them, they are highly likely to be detected.
Even if an adversary avoids detection, their crafted query will be rephrased by the SLM (see Figure 2), which adds the correct signature. [^8] This step is programmatic and cannot be bypassed. As the results in Table 5 demonstrate, an adversary cannot successfully circumvent this security measure even with full knowledge of the bias mechanism.
Method BFT BFT BFT
++ Detection
User Adversary (Bob)
Bias Bob Alice Alice
ChatDoctor* 100.0 100.0 100.0
MedQA 100.0 98.4 100.0
MMLU (med.) 100.0 80.2 91.0
LSE* 100.0 100.0 100.0
LegalBench 100.0 87.6 96.8
MMLU (leg.) 99.8 91.8 98.4
Table 5: Performance of the sudoLLM framework under adversarial conditions This table reports the alignment accuracy under the assumption that an adversary (Bob) has full knowledge of the bias mechanism. “BFT + Detection” column shows the final alignment accuracy using a combined BFT and detection approach. Results demonstrate that the system effectively neutralizes the adversarial threat. “*” indicates the datasets used for BFT training.
@@ 5 Discussions
To illustrate this further, note that if we repeat the attack999 GPT-4o, prefix length 25. in Table 4 with Alice’s BQ instead of Bob’s, our ASR is 93.3%, a huge jump from 0.4% with Bob’s BQ (BFT) and 10.3% with the base model (Inst.). This fact, alongside the other evidence presented here, suggests that (i) the models do recognize the injected bias, and (ii) this hidden bias signal helps the model resolve between the competing interests of language modeling and safety alignment, thus leading to improved alignment and attack performance.